OpenStudy (anonymous):

solve the system of equations 3x + 4y = 5 2x - 3y = -8

OpenStudy (anonymous):

ok, there are two methods u can use. Either Elimination or subsitution, do u know this?

OpenStudy (anonymous):

substitution*

OpenStudy (jiteshmeghwal9):

First find the value of3x or 4y

OpenStudy (anonymous):

no

OpenStudy (jiteshmeghwal9):

Let us find the value of 4y 4y=5-3x

OpenStudy (ujjwal):

3x + 4y = 5......... (1) 2x - 3y = -8........ (2) Multiply (1) by 3 and (2) by 4 and add them. you will get x, substitute the value of x in either equation. you will get the answers.

OpenStudy (jiteshmeghwal9):

Opps! I do it with second method.

OpenStudy (anonymous):

using elimination is better (ujjwal's method)

OpenStudy (jiteshmeghwal9):

Yeah! I am with Ujjawal.